OpenStudy (anonymous):

How should this sentence be expressed in numbers? “A number three times as great as the number a equals four less than A.” 3a = a - 4 3a = 4 - a a + 3 = a - 4 one third a = a - 4

OpenStudy (anonymous):

first

OpenStudy (zepp):

Number equals 4 less than A: a - 4 A number 3 times: 3a 3a = a-4, so first.

OpenStudy (anonymous):

Number is 3 times as great as a , that is 3a and it equals 4 less than a , that is a-4 Hence, 3a=a-4